Balance the chemical equation for the important industrial process: C2H4 + O2 + HCl → C2H4Cl2 + H2O. Given 760.0 kg of molecular oxygen, calculate how many kg of C2H4Cl2 can be produced.

---

These are a mix (maybe) of ionic and covalent compounds: Name these compounds. Spell sulfide with the f not the ph. If the metal is a transition metal, you will need to use roman numerals in parentheses, no space between the metal and the parentheses, e.g. iron(III) oxide Covalent compounds should follow IUPAC naming methods.


What is the name for CS2?  
What is the name for N2O?  
What is the name for Fe2O3?  
What is the name for PBr3?

----

Arrange the following elements in order of increasing electronegativity.

B
Al
In
Ga
O
Answer #1

1) to find killogram of C2H4Cl2 produced

given mass of O2 available for reaction = 760 killogram , mass in gram = mass in killogram x 1000

substitute values to find mass of O2 availble in gram

mass of O2 available for reaction in gram = 760 x 1000 = 760000 gram

molar mass of O2 = 31.9988 gram/mole

no of moles = mass in gram/molar mass , substitute values to find no of moles of O2 available for reaction

no of moles of O2 available for reaction = 760000/31.9988 = 23750.89 moles

balanced chemical reaction is given bellow

2C2H4 + O2 + 4HCl   2C2H4Cl2 + 2H2O   

according to reaction 1 mole of O2 produces 2 moles of C2H4Cl2  

therefore 23750 moles of O2 will produce = 23750.89 x 2 = 47501.78 moles of C2H4Cl2   

no of moles of C2H4Cl2 can be produced = 47501.78 moles , molar mass of C2H4Cl2 = 98.96 gram/moles

mass in gram = no of moles x molar mass , substitute values to find grams of C2H4Cl2 produced

grams of C2H4Cl2 produced = 47501.78 x 98.96 = 4700776.148 gram

mass in killogram = mass in gram/1000 substitute values to find killograms of C2H4Cl2 produced

mass of C2H4Cl2 produced in killogram = 4700776.148/1000 = 4700.776148 killogram

ans 4700.776148 Kg of C2H4Cl2 can be produced from 760 Kg of molecular oxigen

2) to write names of given chemicals

ans name of CS2 is carbon disulfide

name of N2O is dinitrogen monoxide

name of Fe2O3 is Iron(III)oxide   

name of PBr3 is phosphorus tribromide     

3) to find increasing order of electronegativity

increasing order of electronegativity is given bellow

Al,In,Ga,B,O